5. You roll a 12-sided die with equal probability of having the numbers 1 through 12 on the upper side. Your friend covers up the number it shows, but you are given the information that it only has one digit (and therefore can't be 10, 11 and 12). What is the probability it has 5 on the upper side?

6. The crosstabulation below represents the number of households (1,000s) and the household income by the highest level of education for the head of the household. Use the crosstabulation to answer the following questions. a. What is the probability a household is headed by someone with a high school diploma earning $100,000 or more? b. What is the probability a household is headed by someone with a bachelor’s degree earning less than $25,000?

7. 100 high school students are asked if they speak German. 70% of male students answer that they do not speak German, while 40% of female students answer they speak German. It is known that 40% of the responding students are female. If a randomly chosen student answers that he/she speaks German, what is the probability that this student is a female?
Answer #1

We would be looking at Q5 here:

Q5) Given that it is a single digit number, probability that it is 5 is computed here as: (Using Bayes theorem )

P(X = 5 | X <= 9)

= P(X = 5) = P(X <= 9)

= 1 / 9

Therefore 1/9 = 0.1111 is the required probability here.